OpenStudy (anonymous):

Verify the identity. cot ( theta - pi / 2 ) = - tan theta

OpenStudy (anonymous):

I know cos(theta)+sin(1)= Sin(theta)

OpenStudy (anonymous):

Working with Sin(theta-pi/2)

OpenStudy (anonymous):

See enclosed

OpenStudy (anonymous):

I see. I was researching problem further and found this \[\sin(\theta-\pi/2)=\sin(\theta)\cos(\pi/2)-\cos \theta \sin (\pi/2)\]........... Question from where did the \[\cos(\pi/2) \] by itself come from in the second half of the equation.

OpenStudy (anonymous):

This is a standard trigonometric identity sin(A-B)=sin(A)cos(B)-cos(A)sin(B)

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

this was an angle sum and difference identity. I get it now.